Circular No. 5933

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET WILD 3 (1994b) J. V. Scotti and T. Gehrels, Lunar and Planetary Laboratory, report

that they have recovered this comet with the 0.9-m Spacewatch Telescope at Kitt Peak, as shown below. On Feb. 10, there was a coma 12′′ across and tail 0′.29 long in p.a. 292^{o} (m_{2} = 22.3). The indicated correction to the prediction on MPC 18258 (ephemeris on MPC 23139) is ΔT = +0.03 day.

1994 UT  α_{2000}  δ_{2000}  m_{1}
Feb. 10.49233  15^{h}11^{m}48.^{s}72  −4^{o}00′10′′.9
10.49730  15 11 48.96  −4 00 12.0  20.7
10.50515  15 11 49.55  −4 00 12.7  21.2
10.51078  15 11 49.78  −4 00 13.7
10.51564  15 11 50.14  −4 00 15.0
13.50508  15 14 53.30  −4 09 03.6
13.51935  15 14 54.25  −4 09 04.7

A0535+26 R. B. Wilson and B. A. Harmon, Marshall Space Flight Center, NASA;

and M. H. Finger, Computer Sciences Corporation, report for the Compton Gamma Ray Observatory BATSE team: “The outburst reported on IAUC 5931 has continued and intensified through Feb. 8. The daily phaseaveraged pulsed intensity in the band 20-40 keV, relative to the total Crab Nebula flux (pulsed + steady), was 0.2, 0.5, 0.7, 0.8, 0.9, and 1.1 on Feb. 3-8. The daily total intensity (also 20-40 keV), measured using Earth occultation, was 0.4, 0.9, 1.4, 1.9, 2.2, and 2.7 relative to the total Crab Nebula flux. All values stated above are accurate to ± 5 percent. The source is detectable to at least 100 keV.”

SU TAURI B. S. Carter, South African Astronomical Observatory, reports: “Ob-

servation of SU Tau by F. Marang on Jan. 29.89 using the 0.75-m telescope yields magnitudes J = 11.74, H = 9.22, K = 7.14, L = 4.59, which indicates that a dust shell has been ejected.”

1994 February 14 Daniel W. E. Green